# Routoo: Learning to Route to Large Language Models Effectively

### v2

@@ Abstract
Developing foundational large language models (LLMs) is becoming increasingly costly and inefficient. Also, closed-source and larger open-source models generally offer better response quality but come with higher inference costs than smaller models. In this paper, we introduce Routoo, an architecture designed to optimize the selection of LLMs for specific prompts based on performance, cost, and efficiency. Routoo consists of two key components: a performance predictor and a cost-aware decoding. The performance predictor is a lightweight LLM that estimates the performance of various underlying LLMs without needing to execute and evaluate them. The cost-aware decoding then selects the most suitable model based on these predictions and other constraints like cost and latency.
We evaluated Routoo using the MMLU benchmark across 57 domains employing open-source models. Our results show that Routoo matches the performance of the Mixtral 8x7b model while reducing inference costs by one-third. Additionally, by allowing increased costs, Routoo surpasses Mixtral’s accuracy by over 5% at equivalent costs, achieving an accuracy of 75.9%. When integrating GPT4 into our model pool, Routoo nearly matches GPT4’s performance at half the cost and exceeds it with a 25% cost reduction. These outcomes highlight Routoo’s potential to create new SOTA in a cost-effective manner by leveraging the collective knowledge of multiple LLMs.
@@ 1 Introduction
Developing foundational models OpenAI et al. (2024); Touvron et al. (2023); Jiang et al. (2023a, 2024) is capital-intensive, necessitating vast computational resources and extensive, high-quality data Minaee et al. (2024). Furthermore, the field is nearing the upper bounds of network size and data capacity, resulting in progressively marginal enhancements over existing models Udandarao et al. (2024). This scenario echoes a critical juncture in human advancement, where the ’divide and conquer’ methodology emerges as a viable and scalable alternative. This approach entails cultivating domain-specific experts and judiciously harnessing them to forge a composite, high-performance model.
The capabilities of existing LLMs appear to be complementary to a significant degree. An illustrative case is the MMLU benchmark Hendrycks et al. (2021), where selecting the optimal open-source model for each question hypothetically yields an accuracy of 97.5%, at a computational cost akin to a model with 13 billion parameters.  [^1] Further details are provided in Appendix A. In contrast, GPT4 OpenAI et al. (2024) achieves an accuracy of 86.4%, while Mixtral 8x7b Jiang et al. (2024), as the leading open-source model, reaches 70%. These figures suggest considerable scope for integrating LLMs’ knowledge to create new state-of-the-art models. Additionally, it is noteworthy that many practical tasks do not require intricate reasoning and can be efficiently addressed by models of moderate complexity; only a minority of tasks demand the advanced capabilities of a model like GPT4 Zaharia et al. (2024); Ding et al. (2024); Shnitzer et al. (2023); Chen et al. (2023). The landscape of open-source models, with approximately 450,000 entries on Hugging Face, [^2] https://huggingface.co is both dynamic and expansive. The rapid proliferation of new models, particularly smaller, domain-focused ones, poses a challenge in tracking and leveraging the latest advancements effectively.
Figure 1: Routoo architecture has two main components: a performance predictor and a cost-aware decoder. The former predicts the accuracy of experts for a specified query, and the latter chooses the underlying LLM and generates the response by considering the cost and efficiency of each model. The Universe Constructor builds a complementary set of underlying models.
We propose Routoo, an architecture to learn to leverage a universe of trained LLMs automatically and dynamically to create a better-performing model.
@@
The cornerstone of Routoo is a lightweight LLM that predicts the performance of each underlying LLM on queries without the need for actual execution and evaluation. Based on these predictions, a cost-aware decoding algorithm selects the most suitable model, prioritizing factors such as inference cost and speed. For instance, when faced with a task that is predicted to be performed nearly equally well by a 7 billion parameter model or a more extensive 70 billion parameter model, the Routoo will opt for the former when factors like speed and cost-efficiency are prioritized. This approach ensures optimal resource utilization without compromising on quality.
@@
Our architecture marks a significant departure from traditional Mixture of Experts (MoE) models Lieber et al. (2024); Jiang et al. (2024); Fedus et al. (2021); Shazeer et al. (2017). While MoE relies on gating over various expert sub-networks within each layer to predict the next token, it requires all expert parameters to be loaded onto a single, high-end machine. This limitation hinders scalability in the number of experts. In contrast, each ’expert’ within our system operates independently and can be hosted on different machines, potentially utilizing a different neural network architecture. This flexibility means we can incorporate a vast array of experts, ranging from those specializing in system-level Java programming to those adept at curating travel experiences in London. Moreover, Routoo can be tuned to leverage only smaller models, which is particularly beneficial for users without access to high-end computing resources (e.g., GPUs with 100GB). Furthermore, our architecture facilitates easy integration of additional optimization criteria, including cost, speed, and privacy considerations.
@@
We evaluate our Routoo on MMLU benchmark Hendrycks et al. (2021). We show that it achieves competitive performance with Mixtral 8x7b Jiang et al. (2024) while only consuming two-thirds of its inference cost. Increasing the cost budget allows Routoo to outperform Mixtral by 5% at the same cost level, reaching an accuracy of 75.9%. By adding GPT4 OpenAI et al. (2024) as one of the underlying experts, our Routoo achieves competitive performance with GPT4 at half the cost and exceeds it with a 25% cost reduction.
@@
  • We propose Routoo, an LLM-based system designed to intelligently identify the best-performing LLM for a given query while considering constraints (e.g. cost), effectively integrating knowledge of multiple LLMs to create a new state-of-the-art model.
  • We evaluate our architecture on MMLU benchmark, and significantly outperform Mixtral 7x8b by 5% with similar inference cost. Also, our Routoo achieves competitive performance with GPT4 at half the inference cost and surpasses it by reducing the inference cost by 25%.
@@ 2 Related Work
LLM Benchmarks.
In recent years, many benchmarks Wang et al. (2024); Liang et al. (2023); Beeching et al. (2023); Gao et al. (2021); Hendrycks et al. (2021) are created to identify the capability of LLMs. The Massive Multitask Language Understanding  (MMLU) benchmark Hendrycks et al. (2021) contains multiple-choice question-answer (MCQA) pairs from various branches of knowledge. This includes 57 domains, including mathematics, US history, computer science, and law. OpenLLM leaderboard Beeching et al. (2023) includes several question-answering datasets, e.g. HellaSwag Zellers et al. (2019), Truthfulqa Lin et al. (2022), Winogrande Sakaguchi et al. (2019), GSM8k Cobbe et al. (2021), ARC Clark et al. (2018) and MMLU, which is used to compare open-source LLMs. HELM Liang et al. (2023) consists of 42 scenarios covering a variety of use cases. In this work, we utilized the MMLU benchmark Hendrycks et al. (2021) to assess our models. This benchmark spans several domains and is recognized as one of the challenging benchmarks on the OpenLLM leaderboard Lu et al. (2023).
@@
Previous works in selecting the best LLMs mainly focus on identifying the one that generates the most optimal output for a given input. Liu and Liu (2021); Ravaut et al. (2022) proposed specialized scoring or re-ranking models that can be used for the generation tasks (summarisation, here). Speculative decoding Kim et al. (2023); Leviathan et al. (2023) accelerates the decoding of expensive models by using small, efficient decoders for the ’easy’ steps. This approach can complement routing methods. LLM-BLENDER Jiang et al. (2023b) is an ensembling framework to reach better performance by mixing the results of LLMs with a ranking module, followed by a generation module to generate from top candidates of the ranker. FrugalGPT Chen et al. (2023) executes experts sequentially until an expert reaches the acceptable generation performance. HybridLLM Ding et al. (2024) proposes a routing approach to direct queries to the suitable expert, utilizing two underlying LLMs. Different from previous work, our Routoo identifies the most suitable expert without executing the inference of underlying LLMs. Additionally, our approach can be efficiently generalized to scenarios involving many underlying LLMs.
@@ 3.1 Problem Formulation
Given a set of $N$ LLM $e_{1},e_{2},\ldots,e_{N}$ and a set of $m$ queries $q_{1},q_{2},\ldots,q_{m}$, Our goal is to assign the most cost-effective model $e_{i}$ to each query $q_{j}$that can accurately answer the query.
For each model-query pair, we define an evaluation function $\operatorname{eval}(e_{i}(q_{j}))$ [^3] Evaluation function can have the reference as an optional input., which returns a score from 0 to $K-1$ indicating the correctness of the model’s response, where 0 is unacceptable and $K-1$ is optimal. Each response also incurs a cost $\operatorname{cost}(e_{i}(q_{j}))$.
@@
$\max_{\pi}$ $\ \sum_{j=1}^{m}\operatorname{eval}(\pi(q_{j})(q_{j}))$
s.t. $\ \sum_{j=1}^{m}\operatorname{cost}(\pi(q_{j})(q_{j}))\leq B$
Here, $\pi$ is the function that assigns each query to a model, based on both cost and correctness. In this formulation, the cost can be multi-faceted, encompassing aspects like computational cost, speed, etc. However, for simplicity, we consider a singular budget constraint $B$ in this context. To solve this without exhaustively testing every model on every query—a prohibitive approach—we divide the problem into two manageable parts:
@@
    Performance Prediction: This component approximates the evaluation score $\operatorname{eval}(e_{i}(q_{j}))$ for each model on each query without actual execution.
@@
    Cost-Aware Decoder: This step selects the best model for each query, balancing accuracy and cost-effectiveness using the performance predictions from the previous component.
These components work together to dynamically and efficiently allocate models to queries, as illustrated in Figure 1. In the following sections, we will delve deeper into each component.
@@ 3.2 Performance Predictor
Figure 2: The performance predictor: Given the query and an LLM, it outputs the performance prediction $\operatorname{\hat{eval}}(e_{i},q_{j})$.
The performance predictor is a lightweight LLM designed to estimate the effectiveness of each underlying LLM for a given query. It predicts the evaluation score, representing the expected response quality of a model to a query. This prediction is formulated as:
$\operatorname{\hat{eval}}(e_{i},q_{j})=\operatorname{Pred}(e_{i},q_{j}),$ (1)
where $\operatorname{\hat{eval}}(e_{i},q_{j})$ denotes the predicted score of model $e_{i}$ for query $q_{j}$, and $\operatorname{Pred}(e_{i},q_{j})$ is the predictive model.
The process, illustrated in Figure 2, involves:
$\begin{split}\begin{cases}h_{q_{j}}=\operatorname{Enc}(q_{j})\\ h_{e_{i}}=\operatorname{Emb}(e_{i})\\ s_{q_{j},e_{i}}=\operatorname{Linear}(h_{q_{j}}-h_{e_{i}})\end{cases}\end{split}$
where $\operatorname{Enc}(.)$ is the encoder of the input query. Inspired by Radford et al. (2019), we use a decoder-only model as the encoder and extract the embedding of the last token as the representation of the input query. $Emb(.)$ is an embedding layer, where each embedding is assigned to a specific model. $Linear(.)$ is a linear layer to convert the combination of query and model embeddings to the logits of the evaluation score. $s_{q_{i},e_{j}}$ is a list of predicted evaluation scores for a given model and query. The size of this list depends on the number of levels (K) in the evaluation score e.g. 2 for binary evaluation scores.
The model is trained by minimising the cross-entropy loss function Good (1952) of $s_{q_{i},e_{j}}$ and $\operatorname{eval}(e_{j}(q_{i}))$ over all $N$ expert models, and $m$ queries as:
$\operatorname{min}\frac{1}{m}\sum_{i=1}^{m}\sum_{j=1}^{N}\operatorname{CE}(s_{q_{i},e_{j}},\operatorname{eval}(e_{j}(q_{i})))$ (2)
where $\operatorname{CE}(.)$ is the cross-entropy loss.
During the inference time, $\operatorname{argmax}(.)$ is applied for a given model and query as:
$\operatorname{\hat{eval}}(e_{i},q_{j})=\operatorname{argmax}_{(0,...,K-1)}(s_{q_{j},e_{i}})$ (3)
where $\operatorname{argmax}(.)$ is applied over the levels of the evaluation score.
3.3 Cost-Aware Decoder
The second phase of our architecture involves the decoding step, where the estimated scores from the performance predictor are used to determine the optimal assignment of models to queries. Our objective is to maximize the overall effectiveness of the responses within a given budget constraint. The optimization problem is formulated as follows:
$\max_{\pi}$ $\ \sum_{j=1}^{m}\operatorname{\hat{eval}}(\pi(q_{j}),q_{j})$
s.t. $\ \sum_{j=1}^{m}\operatorname{cost}(\pi(q_{j})(q_{j}))\leq B$
where $\pi(.)$ determines the expert assignments and $B$ represents the predefined budget constraint.
We propose a greedy algorithm to approximate this optimization. Noting the challenges of accurately estimating the exact cost of a model for a specific query, instead, we consider the average cost $c_{i}$ for a model $e_{i}$ responding to an average length input and response.
@@ 3.3 Cost-Aware Decoder
while budget < k do
       1. Identify the model $e_{i}^{*}$ that, when added to $U$ forming $U^{*}$, maximizes $S(U^{*})$
       2. Update $U$ to $U^{*}$

### v3

@@ Abstract
LLMs with superior response quality—particularly larger or closed-source models—often come with higher inference costs, making their deployment inefficient and costly. Meanwhile, developing foundational LLMs from scratch is becoming increasingly resource-intensive and impractical for many applications. To address the challenge of balancing quality and cost, we introduce Routoo, an architecture designed to optimize the selection of LLMs for specific prompts based on performance, cost, and efficiency. Routoo provides controllability over the trade-off between inference cost and quality, enabling significant reductions in inference costs for a given quality requirement. Routoo comprises two key components: a performance predictor and cost-aware selector. The performance predictor is a lightweight LLM that estimates the expected performance of various underlying LLMs on a given prompt without executing them. The cost-aware selector module then selects the most suitable model based on these predictions and constraints such as cost and latency, significantly reducing inference costs for the same quality. We evaluated Routoo using the MMLU benchmark across 57 domains employing open-source models. Our results show that Routoo matches the performance of the Mixtral 8x7b model while reducing inference costs by one-third. Additionally, by allowing increased costs, Routoo surpasses Mixtral’s accuracy by over 5% at equivalent costs, achieving an accuracy of 75.9%. When integrating GPT4 into our model pool, Routoo nearly matches GPT4’s performance at half the cost and exceeds it with a 25% cost reduction. These outcomes highlight Routoo’s potential to significantly reduce inference costs without compromising quality, and even to establish new state-of-the-art results by leveraging the collective capabilities of multiple LLMs.
@@ 1 Introduction
LLMs have achieved remarkable success across a wide range of natural language processing tasks. However, this success comes with a significant trade-off between performance and cost: larger models generally offer better response quality but incur higher inference costs than their smaller counterparts. This increased inference cost poses challenges for deploying LLMs in real-world applications where computational resources, latency, and cost-efficiency are critical considerations.
At the same time, developing foundational LLMs  (OpenAI et al., 2024; Dubey et al., 2024; Yang et al., 2024; DeepSeek-AI et al., 2024; Jiang et al., 2024; 2023a; Touvron et al., 2023) from scratch is becoming increasingly capital-intensive, requiring vast computational resources and extensive, high-quality data (Minaee et al., 2024). As the field approaches the limits of network size and data capacity, the improvements gained from training ever-larger models are diminishing (Udandarao et al., 2024). This situation underscores the need for alternative approaches that can achieve high performance without incurring prohibitive development and inference costs.
Moreover, many practical tasks do not necessitate the intricate reasoning capabilities of the largest models like GPT-4; instead, they can be efficiently handled by smaller, more cost-effective models (Zaharia et al., 2024; Ding et al., 2024; Shnitzer et al., 2023; Chen et al., 2023). The capabilities of existing LLMs appear to be complementary to a significant degree. For example, on the MMLU benchmark (Hendrycks et al., 2021), selecting the optimal open-source model for each question could hypothetically yield an accuracy of 97.5% at a computational cost similar to that of a 13-billion-parameter model (see Appendix  A for details). In contrast, GPT-4 (OpenAI et al., 2024) achieves an accuracy of 86.4%, while Mixtral 8×7b (Jiang et al., 2024), one of the leading open-source models, [^1] As of the time of writing this paper. reaches 70%. These observations suggest that integrating the knowledge of multiple LLMs could lead to new state-of-the-art models without the need to train from scratch while significanly reducing the inference cost. However, effectively leveraging the vast and rapidly growing ecosystem of LLMs poses significant challenges. With approximately 450,000 models available on Hugging Face, [^2] https://huggingface.co keeping track of the latest advancements and integrating them efficiently is a non-trivial task. Traditional approaches like Mixture of Experts (MoE) models (Abdin et al., 2024; Lieber et al., 2024; Jiang et al., 2024; Fedus et al., 2021; Shazeer et al., 2017) are limited by the need to load all expert parameters onto a single high-end machine, hindering scalability and flexibility. There is a need for an architecture that can dynamically and efficiently leverage multiple existing LLMs to optimize performance while controlling inference costs.
To address these challenges, we propose Routoo , an architecture designed to optimize the selection of LLMs for specific prompts based on performance, cost, and efficiency. Routoo provides controllability over the trade-off between inference cost and quality, enabling significant reductions in inference costs for a given quality requirement. By intelligently leveraging a universe of trained LLMs, Routoo addresses both the inference cost problem and the development cost of building LLMs from scratch, creating a composite high-performance model without the need for extensive retraining.
@@
Routoo comprises two key components: a performance predictor and cost-aware selector. The performance predictor is a lightweight LLM that estimates the expected performance of various underlying LLMs on a given query without executing them. Based on these predictions and constraints such as cost and latency, the cost-aware selector module selects the most suitable model. For instance, when a task is predicted to be performed nearly equally well by a smaller model or a larger, more expensive model, Routoo will opt for the smaller model when speed and cost-efficiency are prioritized. This approach ensures optimal resource utilization without compromising on quality.
@@
Our architecture marks a significant departure from traditional MoE. While MoE relies on gating over various expert sub-networks within each layer to predict the next token, it requires all expert parameters to be loaded onto a single, high-end machine. This limitation hinders scalability in the number of experts. In contrast, each ’expert’ within our system operates independently and can be hosted on different machines, potentially utilizing a different neural network architecture. This flexibility enables Routoo to incorporate a vast array of experts, ranging from specialized domain models to general-purpose ones, and to scale to a large number of experts without the limitations imposed by MoE models.
@@
We evaluate our Routoo on MMLU benchmark (Hendrycks et al., 2021). We show that it achieves competitive performance with Mixtral 8x7b (Jiang et al., 2024) while only consuming two-thirds of its inference cost. Increasing the cost budget allows Routoo to outperform Mixtral by 5% at the same cost level, reaching an accuracy of 75.9%. By adding GPT4 (OpenAI et al., 2024) as one of the underlying experts, our Routoo achieves competitive performance with GPT4 at half the cost and exceeds it with a 25% cost reduction. These outcomes highlight Routoo’s potential to significantly reduce inference costs without compromising quality, and even to establish new state-of-the-art results by leveraging the collective capabilities of multiple LLMs. By providing controllability over cost and quality trade-offs, Routoo offers an efficient approach to high-performance language modeling without the need for expensive model training from scratch.
@@
  • We propose Routoo, an LLM-based system that intelligently identifies the best-performing LLM for a given query while considering constraints such as cost and latency, effectively integrating the knowledge of multiple LLMs to create a high-performance model without the need for training from scratch.
  • We evaluate our architecture on MMLU benchmark, and significantly outperform Mixtral 8x7b by 5% with similar inference cost. Also, our Routoo achieves competitive performance with GPT4 at half the inference cost and surpasses it by reducing the inference cost by 25%.
@@ 2 Related Work
Mixture-of-Experts.
MoE architecture includes a gating mechanism to integrate various expert sub-networks within each layer, guiding the prediction of the next token (Shazeer et al., 2017). This approach is used in recent foundational models, including Mixtral 8x7b (Jiang et al., 2024). Sukhbaatar et al. (2024) fine-tuned LLaMa 7b (Touvron et al., 2023) on four different domains to create domain-specific expert LLMs. They then proposed combining the FFNs of these expert LLMs to construct an MoE. Tang et al. (2024) proposed merging most parameters while scaling up MLP layrers of Transformer into a weight-ensembling MoE module that dynamically integrates shared and task-specific knowledge based on the input. Wang et al. (2024) introduces a fusion gate at each Transformer layer to generate weights for a weighted average of outputs from a set of pre-trained LoRAs.
Figure 1: Routoo architecture has three main components: a performance predictor, cost-aware selector, and universe constructor. The universe constructor identifies the optimal set of complementary models from all available models. The performance predictor predicts the correctness of experts for a specified query, and the cost-aware selector chooses the underlying model by considering the cost and efficiency of each model.
@@
Previous works in selecting the best LLMs mainly focus on identifying the one that generates the most optimal output for a given input. Liu & Liu (2021); Ravaut et al. (2022) proposed specialized scoring or re-ranking models that can be used for the generation tasks (summarisation, here). Speculative decoding (Kim et al., 2023; Leviathan et al., 2023) accelerates the decoding of expensive models by using small, efficient decoders for the ’easy’ steps. This approach can complement routing methods. LLM-BLENDER (Jiang et al., 2023b) is an ensembling framework to reach better performance by mixing the results of LLMs with a ranking module, followed by a generation module to generate from top candidates of the ranker. FrugalGPT (Chen et al., 2023) executes experts sequentially until an expert reaches the acceptable generation performance. ZOOTER (Lu et al., 2024) proposed a reward-guided routing approach that distills rewards from training queries to train a routing function. HybridLLM (Ding et al., 2024) and RouteLLM (Ong et al., 2024) propose a routing approach to direct queries to the suitable expert, utilizing one strong and one weak LLMs.
Different from previous work, our Routoo identifies the most suitable expert without executing the inference of underlying LLMs. Additionally, our approach can be efficiently generalized to scenarios involving many underlying LLMs (not necessarily a tuple of weak and strong LLMs).
@@ 3.1 Problem Formulation
Given a set of $M$ models $m_{1},m_{2},\ldots,m_{M}$ and a set of $N$ queries $q_{1},q_{2},\ldots,q_{N}$, [^3] Section 3.4 will provide a detailed explanation of how we build the set of models from the pool of all available models. our goal is to assign the most cost-effective model $m_{i}$ to each query $q_{j}$ that can accurately answer the query. Correctness and cost scores of each query-model pair are calculated as:
$s_{m_{i},q_{j}}=\operatorname{Eval}(m_{i}(q_{j}))$
$c_{m_{i},q_{j}}=\operatorname{Cost}(m_{i},q_{j})$
where $\operatorname{Eval}(.)$ computes the correctness score given the generated response $m_{i}(q_{j})$ of model $m_{i}$ for query $q_{j}$. $s_{m_{i},q_{j}}$ is an integer that ranges from 0 to $K-1$, indicating the correctness of the model’s response, where 0 is unacceptable, and $K-1$ is optimal. The cost of executing model $m_{i}$ for query $q_{j}$ is calculated by $\operatorname{Cost}(.)$ function.
@@
$\max_{\pi}$ $\ \sum_{j=1}^{N}s_{\pi(q_{j}),q_{j}}$
s.t. $\ \sum_{j=1}^{N}c_{\pi(q_{j}),q_{j}}\leq B$
Here, $\pi(.)$ is the function that assigns each query to a model from the set of $M$ models, while considering both cost and correctness. In this formulation, the cost can be multi-faceted, encompassing aspects like computational cost, speed, etc. However, for simplicity, we consider a singular budget constraint $B$ in this context. To solve this without exhaustively testing every model on every query—a prohibitive approach—we divide the problem into two manageable parts:
@@
    Performance Predictor. This component approximates the correctness score $s_{m_{i},q_{j}}$ for each model on each query. Given a model $m_{i}$ and a query $q_{j}$, it estimates the model’s performance without generating the response for the query, thereby significantly reducing the inference cost.
@@
    Cost-Aware Selector. This module selects the best model for each query, balancing accuracy and cost-effectiveness by using correctness estimations of performance predictor.
These two components work together to dynamically and efficiently allocate models to queries.
Additionally, given the vast array of LLMs available for text generation, [^4] According to the Huggingface platform (https://huggingface.co), there are nearly 47,000 models available for the text generation task. and considering the practical constraints on the number of models that can be actively served, we introduce Universe Constructor, which builds an initial complementary universe of models ($m_{1},m_{2},\ldots,m_{M}$) from a set of all available models. This universe aims to maximize performance by ensuring that the selected models are complementary to each other.
All three components are illustrated in Figure 1. In the following sections, we will delve deeper into each component.
@@ 3.2 Performance Predictor
The performance predictor is a lightweight LLM designed to estimate the effectiveness of each underlying LLM for a given query. It estimates the output of evaluation function ($s_{m_{i},q_{j}}$, introduced in Section 3.1) without executing model $m_{i}$ on query $q_{j}$. This prediction is formulated as:
$\hat{s}_{m_{i},q_{j}}=\operatorname{Pred}(m_{i},q_{j})$
@@
while budget $<$ M do
       1. Identify the model $m_{i}^{*}$ that when added to $U$ forming $U^{*}$ which maximizes $S(U^{*})$
       2. Update $U$ by adding $m_{i}^{*}$